OpenStudy (anonymous):

The numbers 20,21,29 form a Pythagorean triple. Which other set of numbers is a Pythagorean triple? 21,22,30 40,42,58 60,66,87 70,84,116

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Pythagorean_triples

OpenStudy (anonymous):

second one

OpenStudy (anonymous):

40,42,58 (58)^2=(40)^2+(42)^2 3364=1600+1764 3364=3364